OpenStudy (anonymous):

f(x)=-1/(x+1)^2 find the vertical asymptote horizontal asymtote slanted asymtote x int y int holes domain and range

OpenStudy (anonymous):

please don't make it complicated, its late and i need sleep

OpenStudy (anonymous):

never mind! instead do it for the equation 2x^2-x-1/x^2-4x=3

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (coconutjj):

Vertical Asymptote: Factor out denominator... x(x-4)

OpenStudy (jhannybean):

Vertical asymptote: \((x+1)^2=0\)

OpenStudy (anonymous):

on my calculator it says the va is 3?

OpenStudy (jhannybean):

No.

OpenStudy (coconutjj):

Yes, 3 is ONE of the V A's

OpenStudy (jhannybean):

\[(x+1)^2=0\]Solve for x by square rooting both sides. Then isolate the x.

OpenStudy (coconutjj):

Oh But he said this: "never mind! instead do it for the equation 2x^2-x-1/x^2-4x=3"

OpenStudy (jhannybean):

Oh, didn't see that :P

OpenStudy (anonymous):

oh ha sorry @Jhannybean

OpenStudy (anonymous):

i have the asymptotes what about the rest?

OpenStudy (anonymous):

especialy x's y's and hole

OpenStudy (anonymous):

can someone just tell me the equations for each part, it would help a lot!

OpenStudy (jhannybean):

\[y=\frac{2x^2-x-1}{x^2-4x-3}\]\[\begin{align} 2x^2-x-1 \\&= x^2-x-2 \\&=(x-2)(x+1) \\&=\left(x-\frac{2}{2}\right)\left(x+\frac{1}{2}\right) \\&=(x-1)(2x+1) \end{align}\] \[y=\frac{(x-1)(2x+1)}{(x-3)(x-1)}\]

OpenStudy (jhannybean):

Well,... I would set it up in that fashion to find everything that we need.

OpenStudy (jhannybean):

vertical asymptotes: just find the 0's in the denominator: \((x-3)(x-1)\) horizontal asymptotes: compare \(\dfrac{2x^2}{x^2}\). Since the degree of the polynomials are the same, the HA = \(\dfrac{2}{1} = 2\) No Slant.

OpenStudy (anonymous):

awesome thanks

OpenStudy (jhannybean):

For x and y intercepts: x-int: \(0=\dfrac{(x-1)(2x+1)}{(x-3)(x-1)}\) . Cross multiply, and solve for x-values. y-int: \(y=\dfrac{2(0)^2-(0)-1}{(0)^2-4(0)-3}\) Solve

OpenStudy (jhannybean):

Holes: Happen when you have a removable factor from your function:\[y=\frac{\color{red}{(x-1)}(2x+1)}{\color{red}{(x-1)}(x-3)}\] what value of x would make the denominator = 0?

OpenStudy (jhannybean):

Domain: \[y=\frac{(x-1)(2x+1)}{(x-3)(x-1)}\]Which values of x would cause the denominator to equal 0? solve for \((x-3)(x-1)=0\) Range: find the range by graphing your function.